# Controlling the Flow: Stability and Convergence for Stochastic Gradient Descent with Decaying Regularization

### v1

@@ 2.1 General convergence results
First, we present a general convergence statement for reg-SGD to the minimum-norm solution, both in the almost sure sense as well as the $L^{2}$-sense. The assumptions on the sequence of step-sizes $\alpha$ are similar to the Robbins-Monro step-size conditions. Regarding the sequence of regularization parameters $\lambda$, the assumptions for deriving almost sure convergence to the minimum-norm solution reflect the competing goals of using the strong convexity of $f_{\lambda}$ for $\lambda>0$ and having sufficiently fast convergence of $x_{\lambda}\to x_{\ast}$. Compared to the almost sure convergence statement in Theorem 2.1, the second result, Theorem 2.2, establishes convergence in $L^{2}$ under arguably much weaker assumptions. In particular, the sequence $\lambda$ is allowed to decay at a very slow rate and no prior knowledge of the rate of convergence for $x_{\lambda}\to x_{\ast}$ is required.
We stress that we do not impose any boundedness assumptions of the reg-SGD scheme. In particular, the fact that $\sup_{k\in\mathbb{N}_{0}}X_{k}<\infty$ almost surely is a consequence of Theorem 2.1 which is guaranteed by the retracting force of the Tikhonov regularization.
@@
Suppose that 1.1 and 1.2 are fulfilled and let $(X_{k})_{k\in\mathbb{N}_{0}}$ be generated by (4) with predictable (random) step-sizes and regularization parameters that are uniformly bounded from above. Moreover, we assume that almost surely the sequence $\lambda$ is decreasing to 00 and that
@@
One can question how to verify the third assumption in (6) for practical applications. In Appendix D, we quantify the distance between $x_{\lambda}$ and $x_{\ast}$ in linear inverse problems satisfying a source condition, as well as in the situation, where $f$ satisfies a Łojasiewicz inequality. In general, one has no control for $\|x_{\ast}\|_{\mathcal{X}}^{2}-\|x_{\lambda}\|_{\mathcal{X}}^{2}$, see torralba1996convergence . We thus present a second result on $L^{2}$-convergence that holds also under a simpler condition. Here we require deterministic step-sizes and regularization parameters.
@@
Suppose that 1.1 and 1.2 are fulfilled and let $(X_{k})_{k\in\mathbb{N}_{0}}$ be generated by (4) with decreasing $(\lambda_{k})_{k\in\mathbb{N}}$. Moreover, assume that $\lambda_{k}\to 0$ and (6), or, alternatively, that
@@ 2.2 Convergence rates
Note that, due to Theorem 2.2, one has $\mathbb{E}[\|X_{k}-x_{\ast}\|_{\mathcal{X}}^{2}]\to 0$ if $q>p$ and $p+q<1$. However, we can further derive the following convergence rates.
@@
Suppose that 1.1 and 1.2 are satisfied. Let $C_{\alpha},C_{\lambda}>0$, $p\in(0,\frac{1}{2}]$ and $q\in(p,1-p]$. Let $(X_{k})_{k\in\mathbb{N}_{0}}$ be generated by (4) with $\alpha_{k}=C_{\alpha}k^{-q}$ and $\lambda_{k}=C_{\lambda}k^{-p}$. If $q=1-p$ we additionally assume that $2C_{\lambda}C_{\alpha}>1-q$. Then it holds that $\lim_{k\to\infty}\mathbb{E}[\|X_{k}-x_{\ast}\|_{\mathcal{X}}^{2}]=0$ and
@@
Finally, we derive almost sure convergence rates for regularized SGD. We highlight that Theorem 2.4 additionally gives almost sure convergence of reg-SGD to the minimum-norm solution for a specific choice of schedules without additional assumptions on the rate of convergence for $\|x_{\ast}\|-\|x_{\lambda}\|$.
@@
Suppose that 1.1 and 1.2 are satisfied. Let $C_{\alpha},C_{\lambda}>0$, $p\in(0,\frac{1}{3})$ and $q\in(\frac{p+1}{2},1-p)$. Let $(X_{k})_{k\in\mathbb{N}_{0}}$ be generated by (4) with $\alpha_{k}=C_{\alpha}k^{-q}$ and $\lambda_{k}=C_{\lambda}k^{-p}$. Then, for any $\beta\in(0,2q-1)$, it holds that $\lim_{k\to\infty}\|X_{k}-x_{\ast}\|_{\mathcal{X}}=0$ almost surely and
@@
In Theorem C.5 of the appendix we also provide a theorem on convergence rates for deterministic reg-GD (3) with polynomial step-size and regularization schedules.
@@ 2.3 Refinements under Łojasiewicz condition
with $\xi=\frac{1-\tau}{2}$, see maulen2024tikhonov . We provide further discussion in Appendix D. Note that (8) is sufficient to guarantee (9), however, in the subsequent convergence rates we rely only on (9). Now we use that
@@
so that we can bound the distance to the minimum-norm solution by (9) and the statements derived in Theorem 2.3 and Theorem 2.4.
Regarding the convergence in $L^{2}$, Theorem 2.3 together with (9) implies the following strong convergence rates in $L^{2}$:
@@
Suppose that the conditions of Theorem 2.3 are satisfied and assume that (9) is in place for some $\xi>0$. Then it holds that
@@
For almost sure convergence, Theorem 2.4 together with (9) implies strong a.s. convergence rates:
@@
Suppose that the conditions of Theorem 2.4 are satisfied and assume that (9) is in place for some $\xi>0$. Then for all $\beta\in(0,2q-1)$ it holds that
$\|X_{k}-x_{\ast}\|_{\mathcal{X}}^{2}=\mathcal{O}(k^{-\min(1-q-p,\beta-p,2\xi p)})\quad\text{ almost surely.}$.
Let $\varepsilon>0$ and choose $\beta=2q-1-\varepsilon$. Then, for the optimal values $q=\frac{2}{3}$ and $p=\frac{1}{6\xi+3}$ we get
$\|X_{k}-x_{\ast}\|_{\mathcal{X}}^{2}=\mathcal{O}(k^{-\frac{2\xi}{6\xi+3}-\varepsilon}),\quad\text{ almost surely.}$.
In Figure 2, we illustrate the convergence rate of $\|X_{k}-x_{\ast}\|_{\mathcal{X}}^{2}$ depending on the decay-rates $p,q$ of schedules $\alpha$ and $\lambda$ in the situation where $f$ satisfies a Polyak-Łojasiewicz inequality, i.e. (8) is satisfied with $\tau=\frac{1}{2}$ and, thus, (9) is satisfied with $\xi=\frac{1}{4}$. In Section A.2 we provide a numerical experiment studying the behavior of convergence when implementing reg-SGD for different choices of $\alpha$ and $\lambda$.
Figure 2: Optimal choices of $p$ and $q$. Left: convergence rate for $\mathbb{E}[\|X_{k}-x_{\ast}\|_{\mathcal{X}}^{2}]$ in the situation of Corollary 2.6. Right: almost sure convergence rate for $\|X_{k}-x_{\ast}\|_{\mathcal{X}}^{2}$ in the situation of Corollary 2.7 under the Polyak-Łojasiewicz inequality.
@@ 3.1 Why is reg-SGD important?
$x_{\ast}:=\operatorname{arg\min}\Big{\{}\|\hat{x}\|_{\mathcal{X}}\Big{|}\hat{x}\in\operatorname{arg\min}_{x\in\mathcal{X}}\|Ax-y\|_{\mathcal{Y}}\Big{\}}\,.$
@@
In what follows, we present results from an experiment on a task of image reconstruction based on the Radon transformation. This experiment demonstrates the relevance of carefully tuning decreasing step-size and regularization schedules. Two additional experiments are provided in Appendix A that highlight the performance of our theoretically derived optimal schedules.
@@ 3.2 Fine-tuned reg-SGD for X-ray tomography
[…]
@@
We carried out an experiment, reconstructing an image from it’s Radon transform, solving the ill-posed optimization problem using SGD and reg-SGD with our optimal step-size schedule and a more aggressive regularization schedule. All details of the implementation are provided in Section A.1. The experiment demonstrates the strength of our fine-tuned step-size and regularization schedules. While our optimal schedules ($p=\frac{1}{3}$, $q=\frac{2}{3}$) yield fast convergence to the minimum-norm solution, a more aggressive schedule ($p=q=\frac{2}{3}$) stagnates at a suboptimal level. More critically, vanilla SGD with theoretically optimal step-sizes even fails to produce feasible reconstructions. To illustrate this, in Figure 4 we compare the reconstructed images from reg-SGD with the optimal rates from our analysis, reg-SGD with more aggressive rates, vanilla SGD, and the minimum-norm solution $x_{\ast}$, which is computed via the Moore-Penrose pseudoinverse $x_{\ast}=A^{\dagger}y$. Additionally, we plot both the expected and a.s. optimality gap in Figure 5 as well as the $L^{2}$- and pathwise-error to the minimum-norm solution Figure 5. In this experiment, SGD shows faster convergence in terms of the optimality gap, but ultimately fails to converge to the minimum-norm solution.
@@ 4 Conclusion and future work
We analyzed convergence properties of SGD with decreasing Tikhonov regularization. For convex optimization problems that may have infinitely many solutions, we showed that the regularization can always be chosen to guarantee convergence (almost surely and in $L^{2}$) to the minimum-norm solution. In fact, we provided guidance on explicit choices for polynomial step-size and regularization schedules that ensure best (in the sense of our upper bounds) convergence rates. On the way we revealed interesting mathematical insight into the effect of regularization. In contrast to plain vanilla SGD, boundedness of the approximation sequence is always ensured. A number of concrete applications was provided to show that our theoretical best schedules indeed are consistent with experimental observations, specifically in the experiments of Section A.2. Since our analysis is limited to the smooth convex setting without constraints, for future work it could be interesting to
  • extend results beyond the convexity assumption on $f$, e.g. using gradient domination properties fatkhullin2022 ,
@@
  • use decreasing regularization schedules to better understand the relation of implicit and explicit regularization present in SGD,
@@
Acknowledgment
The authors would like to thank Adrian Riekert for helpful discussions at the beginning of this project.
@@ Appendix A Omitted details and additional numerical experiments
In the following section, we give a detailed description of the implementation for our numerical experiment conducted in Section 3. Moreover, we provide two additional experiments.
@@ A.1 Implementation details of the Radon transform
In the example of Section 3.2, we discretize the Radon transform using a fixed set of $32$ equally spaced projection angles $\theta_{i}\in[0,\pi)$, $i=1,\dots,32$ and use $100$ parallel rays per angle. The unknown image is defined on a $128\times 128$ pixel grid and represented as a vector $x^{\dagger}\in\mathbb{R}^{d}\cong\mathbb{R}^{128\times 128}$, $d=128^{2}$. Given any image $x\in\mathbb{R}^{d}$ its discretized Radon transform is implemented as a matrix-vector product $Ax$, where $A\in\mathbb{R}^{K\times d}$ is the forward operator, and $K=32\times 100$, corresponds to the total number of measurements (i.e., the number of angle-ray combinations). Each row of $A$ represents a discrete line integral along one ray at a given projection angle. The objective function is then defined as
@@
We implemented both SGD and reg-SGD by partitioning the forward operator $A\in\mathbb{R}^{K\times d}$ into blocks $A_{i}\in\mathbb{R}^{100\times d}$, each corresponding to a fixed projection angle $\theta_{i}$, $i=1,\dots,32$. At each iteration, the angle $\theta_{i}$ is sampled uniformly at random, and the gradient of $f$ is approximated by $\nabla f_{i}(x)=A_{i}^{\top}(A_{i}x-g_{\theta_{i}})\in\mathbb{R}^{d}$ and additionally perturbed by independent noise following a multivariate normal distribution with zero mean and covariance $0.5^{2}\cdot{\mathrm{Id}}$. For SGD we chose the learning rate schedule $\alpha_{k}=20k^{-1/2}$. For reg-SGD we chose $\alpha_{k}=20k^{-2/3}$ and regularization $\lambda_{k}=0.01k^{-1/3}$. Moreover, we compare to reg-SGD with $\alpha_{k}=20k^{-2/3}$ and regularization $\lambda_{k}=0.01k^{-2/3}$, i.e.,reg-SGD with a too fast decay of regularization. We initialize all algorithms for each repetition at zero.

### v2

@@ 2.1 General convergence results
First, we present a general convergence statement for reg-SGD to the minimum-norm solution, both in the almost sure sense as well as the $L^{2}$-sense. The assumptions on the sequence of step-sizes $\alpha$ are similar to the Robbins-Monro step-size conditions. Regarding the sequence of regularization parameters $\lambda$, the assumptions for deriving almost sure convergence to the minimum-norm solution reflect the competing goals of using the strong convexity of $f_{\lambda}$ for $\lambda>0$ and having sufficiently fast convergence of $x_{\lambda}\to x_{\ast}$. Compared to the almost sure convergence statement in Theorem˜2.1, the second result, Theorem˜2.2, establishes convergence in $L^{2}$ under arguably much weaker assumptions. In particular, the sequence $\lambda$ is allowed to decay at a very slow rate and no prior knowledge of the rate of convergence for $x_{\lambda}\to x_{\ast}$ is required.
We stress that we do not impose any boundedness assumptions of the reg-SGD scheme. In particular, the fact that $\sup_{k\in\mathbb{N}_{0}}X_{k}<\infty$ almost surely is a consequence of Theorem˜2.1 which is guaranteed by the retracting force of the Tikhonov regularization.
@@
Suppose that Section˜1.1 and Section˜1.1 are fulfilled and let $(X_{k})_{k\in\mathbb{N}_{0}}$ be generated by (4) with predictable (random) step-sizes and regularization parameters that are uniformly bounded from above. Moreover, we assume that almost surely the sequence $\lambda$ is decreasing to 0 and that
@@
One can question how to verify the third assumption in (7) for practical applications. In Appendix˜E, we quantify the distance between $x_{\lambda}$ and $x_{\ast}$ in linear inverse problems satisfying a source condition, as well as in the situation, where $f$ satisfies a Łojasiewicz inequality. In general, one has no control for $\|x_{\ast}\|_{\mathcal{X}}^{2}-\|x_{\lambda}\|_{\mathcal{X}}^{2}$, see torralba1996convergence . We thus present a second result on $L^{2}$-convergence that holds also under a simpler condition. Here we require deterministic step-sizes and regularization parameters. Our requirements in (8) are very similar to the ones needed in the deterministic setting (bakushinsky1994illposed, , Theorem 5.1 and Theorem 5.2) and are motivated by the corresponding deterministic result in continuous time (cominetti2008strong, , Theorem 2.2).
@@
Suppose that Section˜1.1 and Section˜1.1 are fulfilled and let $(X_{k})_{k\in\mathbb{N}_{0}}$ be generated by (4) with deterministic step-sizes and deterministic and decreasing regularization parameters $(\lambda_{k})_{k\in\mathbb{N}}$. Moreover, assume that $\lambda_{k}\to 0$ and (7), or, alternatively, that
@@ 2.2 Convergence rates
Note that, due to Theorem˜2.2, one has $\mathbb{E}[\|X_{k}-x_{\ast}\|_{\mathcal{X}}^{2}]\to 0$ if $q>p$ and $p+q<1$. However, we can further derive the following convergence rates.
@@
Suppose that Section˜1.1 and Section˜1.1 are satisfied. Let $C_{\alpha},C_{\lambda}>0$, $p\in(0,\frac{1}{2}]$ and $q\in(p,1-p]$. Let $(X_{k})_{k\in\mathbb{N}_{0}}$ be generated by (4) with $\alpha_{k}=C_{\alpha}k^{-q}$ and $\lambda_{k}=C_{\lambda}k^{-p}$. If q=1pq=1-p we additionally assume that $2C_{\lambda}C_{\alpha}>1-q$. Then it holds that $\lim_{k\to\infty}\mathbb{E}[\|X_{k}-x_{\ast}\|_{\mathcal{X}}^{2}]=0$ and
@@
Finally, we derive almost sure convergence rates for regularized SGD. We highlight that Theorem˜2.4 additionally gives almost sure convergence of reg-SGD to the minimum-norm solution for a specific choice of schedules without additional assumptions on the rate of convergence for $\|x_{\ast}\|-\|x_{\lambda}\|$.
@@
Suppose that Section˜1.1 and Section˜1.1 are satisfied. Let $C_{\alpha},C_{\lambda}>0$, $p\in(0,\frac{1}{3})$ and $q\in(\frac{p+1}{2},1-p)$. Let $(X_{k})_{k\in\mathbb{N}_{0}}$ be generated by (4) with $\alpha_{k}=C_{\alpha}k^{-q}$ and $\lambda_{k}=C_{\lambda}k^{-p}$. Then, it holds that $\lim_{k\to\infty}\|X_{k}-x_{\ast}\|_{\mathcal{X}}=0$ almost surely and for any $\beta\in(0,2q-1)$
@@
In Theorem˜D.5 of the appendix we also provide a theorem on convergence rates for deterministic reg-GD (3) with polynomial step-size and regularization schedules.
@@ 2.3 Refinements under Łojasiewicz condition
with $\xi=\frac{1-\tau}{2}$, see maulen2024tikhonov . We provide further discussion in Appendix˜E. Note that (9) is sufficient to guarantee (10), however, in the subsequent convergence rates we rely only on (10). Now we use that
@@
so that we can bound the distance to the minimum-norm solution by (10) and the statements derived in Theorem˜2.3 and Theorem˜2.4.
Regarding the convergence in $L^{2}$, Theorem˜2.3 together with (10) implies the following strong convergence rates in $L^{2}$:
@@
Suppose that the conditions of Theorem˜2.3 are satisfied and assume that (10) is in place for some $\xi>0$. Then it holds that
@@
For almost sure convergence, Theorem˜2.4 together with (10) implies strong a.s. convergence rates:
@@
Suppose that the conditions of Theorem˜2.4 are satisfied and assume that (10) is in place for some $\xi>0$. Then for all $\beta\in(0,2q-1)$ it holds that
$\|X_{k}-x_{\ast}\|_{\mathcal{X}}^{2}=\mathcal{O}(k^{-\min(1-q-p,\beta-p,2\xi p)})\quad\text{ almost surely.}$
Let $\varepsilon>0$ and choose $\beta=2q-1-\varepsilon$. Then, for the optimal values $p=\frac{1}{6\xi+3}$ and $q=\frac{2}{3}$ we get
$\|X_{k}-x_{\ast}\|_{\mathcal{X}}^{2}=\mathcal{O}(k^{-\frac{2\xi}{6\xi+3}-\varepsilon}),\quad\text{ almost surely.}$
In Figure˜2, we illustrate the convergence rate of $\|X_{k}-x_{\ast}\|_{\mathcal{X}}^{2}$ depending on the decay-rates $p,q$ of schedules $\alpha$ and $\lambda$ in the situation where $f$ satisfies a Polyak-Łojasiewicz inequality, i.e. (9) is satisfied with $\tau=\frac{1}{2}$ and, thus, (10) is satisfied with $\xi=\frac{1}{4}$. In Section˜A.2 we provide a numerical experiment studying the behavior of convergence when implementing reg-SGD for different choices of $\alpha$ and $\lambda$.
Figure 2: Optimal choices of $p$ and $q$. Left: convergence rate for $\mathbb{E}[\|X_{k}-x_{\ast}\|_{\mathcal{X}}^{2}]$ in the situation of Section˜2.3. Right: almost sure convergence rate for $\|X_{k}-x_{\ast}\|_{\mathcal{X}}^{2}$ in the situation of Section˜2.3 under the Polyak-Łojasiewicz inequality.
@@ 3.1 Why is reg-SGD important?
$x_{\ast}:=\operatorname*{arg\,min}\Big\{\|\hat{x}\|_{\mathcal{X}}\Big|\hat{x}\in\operatorname*{arg\,min}_{x\in\mathcal{X}}\|Ax-y\|_{\mathcal{Y}}\Big\}\,.$
@@
In what follows, we present results from an experiment on a task of image reconstruction based on the Radon transformation. This experiment demonstrates the relevance of carefully tuning decreasing step-size and regularization schedules. Two additional experiments are provided in Appendix˜A that highlight the performance of our theoretically derived optimal schedules.
@@ 3.2 Fine-tuned reg-SGD for X-ray tomography
$f\mapsto R_{\theta}[x](t)=\int_{\mathbb{R}}x(t\cos(\theta)-s\sin(\theta),t\sin(\theta)+s\cos(\theta))\,ds$
@@
We carried out an experiment, reconstructing an image from it’s Radon transform (see Figure˜3) solving the ill-posed optimization problem using SGD and reg-SGD with our optimal step-size schedule and a more aggressive regularization schedule. All details of the implementation are provided in Section˜A.1. The experiment demonstrates the strength of our fine-tuned step-size and regularization schedules. While our optimal schedules ($p=\frac{1}{3}$, $q=\frac{2}{3}$) yield fast convergence to the minimum-norm solution, a more aggressive schedule ($p=q=\frac{2}{3}$) stagnates at a suboptimal level. More critically, vanilla SGD with theoretically optimal step-sizes even fails to produce feasible reconstructions. To illustrate this, in Figure˜4 we compare the reconstructed images from reg-SGD with the optimal rates from our analysis, reg-SGD with more aggressive rates, vanilla SGD, and the minimum-norm solution $x_{\ast}$, which is computed via the Moore-Penrose pseudoinverse $x_{\ast}=A^{\dagger}y$. Additionally, we plot both the expected and a.s. optimality gap in Figure˜5 as well as the $L^{2}$- and pathwise-error to the minimum-norm solution Figure˜5. In this experiment, SGD shows faster convergence in terms of the optimality gap, but ultimately fails to converge to the minimum-norm solution.
@@ 4 Conclusion and future work
We analyzed convergence properties of SGD with decreasing Tikhonov regularization. For convex optimization problems that may have infinitely many solutions, we showed that the regularization can always be chosen to guarantee convergence (almost surely and in $L^{2}$) to the minimum-norm solution. In fact, we provided guidance on explicit choices for polynomial step-size and regularization schedules that ensure best (in the sense of our upper bounds) convergence rates. On the way we revealed interesting mathematical insight into the effect of regularization. In contrast to plain vanilla SGD, boundedness of the approximation sequence is always ensured. A number of concrete applications was provided to show that our theoretical best schedules indeed are consistent with experimental observations, specifically in the experiments of Section˜A.2. Since our analysis is limited to the smooth convex setting without constraints, for future work it could be interesting to
  • extend results beyond the convexity assumption on $f$, e.g. using gradient domination properties fatkhullin2022 or the tangential cone condition which is commonly employed in iterative regularization methods for non-linear inverse problems KaltenbacherNeubauerScherzer2008 ,
@@
  • use decreasing regularization schedules to better understand the relation of implicit and explicit regularization present in SGD, and
@@
Acknowledgments
We thank the reviewers and the area chair for their valuable feedback. Our sincere thanks to Adrian Riekert, whose discussions and feedback were instrumental in shaping the project from its outset. SK acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – CRC/TRR 388 ”Rough Analysis, Stochastic Dynamics and Related Fields” – Project ID 516748464.
@@ Appendix A Omitted details and additional numerical experiments
In the following section, we give a detailed description of the implementation for our numerical experiment conducted in Section˜3. Moreover, we provide two additional experiments.
@@ A.1 Implementation details of the Radon transform
In the example of Section˜3.2, we discretize the Radon transform using a fixed set of $32$ equally spaced projection angles $\theta_{i}\in[0,\pi)$, $i=1,\dots,32$ and use $100$ parallel rays per angle. The unknown image is defined on a $128\times 128$ pixel grid and represented as a vector $x^{\dagger}\in\mathbb{R}^{d}\cong\mathbb{R}^{128\times 128}$, $d=128^{2}$. Given any image $x\in\mathbb{R}^{d}$ its discretized Radon transform is implemented as a matrix-vector product $Ax$, where $A\in\mathbb{R}^{K\times d}$ is the forward operator, and $K=32\times 100$, corresponds to the total number of measurements (i.e., the number of angle-ray combinations). Each row of $A$ represents a discrete line integral along one ray at a given projection angle. The objective function is then defined as
@@
We implemented both SGD and reg-SGD by partitioning the forward operator $A\in\mathbb{R}^{K\times d}$ into blocks $A_{i}\in\mathbb{R}^{100\times d}$, each corresponding to a fixed projection angle $\theta_{i}$, $i=1,\dots,32$. At each iteration, the angle $\theta_{i}$ is sampled uniformly at random, and the gradient of $f$ is approximated by $\nabla f_{i}(x)=A_{i}^{\top}(A_{i}x-g_{\theta_{i}})\in\mathbb{R}^{d}$ and additionally perturbed by independent noise following a multivariate normal distribution with zero mean and covariance $0.5^{2}\cdot{\mathrm{Id}}$. For SGD we chose the step-size schedule $\alpha_{k}=20k^{-1/2}$. For reg-SGD we chose $\alpha_{k}=20k^{-2/3}$ and regularization $\lambda_{k}=0.01k^{-1/3}$. Moreover, we compare to reg-SGD with $\alpha_{k}=20k^{-2/3}$ and regularization $\lambda_{k}=0.01k^{-2/3}$, i.e., reg-SGD with a too fast decay of regularization. We initialize all algorithms for each repetition at zero.